# From Coarse to Fine-Grained Emotion Annotation: An Immediate Recall Paradigm with Validation through Physiological Evidence and Recognition Performance

### v1

@@ IV-B Autonomic Response Analysis
GSR serves as a sensitive indicator of sympathetic nervous system activity and emotional arousal [47], with SCR being key markers of autonomic activation [48]. We analyzed the occurrence and characteristics of SCRs within the 4-second event windows surrounding the annotated emotional events, comparing these responses to the pre-trial 10-second baseline period. Typically, during the baseline period, the GSR signal often exhibited a gradually declining trend, indicative of a resting or relaxing state.
To illustrate the differential autonomic response patterns across emotional categories, we present representative physiological recordings for emotions with contrasting arousal levels. Fig. 10 displays GSR, PPG, and ECG signals from a fear trial, exemplifying high-arousal emotional responses. In marked contrast to the baseline’s declining GSR trajectory, fear-annotated events consistently triggered distinct SCRs—rapid increases in skin conductance amplitude characteristic of sympathetic activation [49]. Notably, Events 3 and 11 demonstrate pronounced SCRs immediately following the annotated timestamps, while Event 1 exhibits a sustained conductance elevation. Event 7 shows a more moderate SCR, potentially reflecting individual variability in fear intensity perception, yet still deviating significantly from baseline patterns.continuously declining baseline.
Figure 10: The PPG, GSR and ECG signals recorded for a complete trial to the emotion of fear. The upper part of the figure shows the overall change trend of these physiological signals during the experiment: the initial gray background area is the baseline, and the red vertical dotted line is the marked timestamp. The lower part of the figure zooms in on events 1, 3, 7 and 11, showing the details of the physiological signals around the emotional start time recalled by the participant.
In stark contrast, Fig. 11 presents physiological recordings from a sadness trial, representing low-arousal emotional processing. Throughout both baseline and stimulus presentation periods, the GSR signal maintained its characteristic declining trajectory without significant perturbations. Critically, the annotated sadness events failed to elicit discernible SCRs, the GSR signal continued its downward trend uninterrupted. This absence of phasic sympathetic responses during participant-reported emotional moments provides compelling evidence for the arousal-specific nature of autonomic activation patterns.
Figure 11: The PPG, GSR, and ECG signals recorded during a complete experiment for the emotion sadness. The figure shows the overall change trend of these physiological signals during the experiment: the initial gray background area is the baseline, and the red vertical dotted line is the marked timestamp.
@@
Fig. 12(a) presents the mean values of this SCR-based metric for each of the six basic emotion categorically, ’Fear’ events exhibited the highest GSR increase percentage at 98.8% , followed by ’Surprise’ 92.6%, ’Happiness’ 87.2%, , and ’Anger’ 85.2%, distinctly grouped by high and low arousal levels. As evident from the figure, high-arousal emotions consistently elicited stronger or more frequent SCRs. Conversely, low-arousal emotions ’Sadness’ 7.4%, and ’Disgust’ 4.6% showed substantially lower mean values for this metric, indicating a significantly attenuated phasic sympathetic response. This dichotomy is further emphasized in Fig. 12(b), which directly compares the mean GSR Increase Percentage between the aggregated high-arousal and low-arousal emotion groups. The high-arousal group demonstrated a mean GSR increase percentage of 91.0% , whereas the low-arousal group showed a mean of only 6.0%.
@@ IV-C Comparative of Immediate versus Delayed Recall Annotation Paradigms
To evaluate the effectiveness of the proposed annotation approach, we conducted a controlled comparative experiment examining the differences between immediate and delayed recall annotation methods. The study employed a within-participant design with the same 20 participants from the FIRMED. We selected 48 videos from the SEED-VII dataset as stimuli, comprising 8 videos for each of the six basic emotions. The experimental protocol was carefully balanced: the first 24 videos utilized the immediate recall annotation paradigm (as in FIRMED), while the remaining 24 videos employed the delayed recall approach (as in SEED-VII). This design allowed us to assess the impact of memory bias on annotation quality through both subjective and objective measures.
@@
First, participants rated their subjective confidence regarding annotation accuracy using a 7-point Likert scale (1 = completely unsure, 7 = extremely certain) [50]. Descriptive statistics indicated that the mean confidence in the immediate annotation condition (Mean = 6.85, SD = 0.21) was numerically higher than in the delayed condition (Mean = 5.91, SD = 1.35). Crucially, there was a significant difference in response variability between the conditions. A Levene’s test [51] was conducted to statistically evaluate the equality of variances, revealing significant heterogeneity (F(1, 38) = 47.2, $p$ <0.001). To further quantify this relative dispersion, we computed the coefficient of variation (CV) [52], which calculated as CV = SD/Mean. The CV for the immediate condition was only 3.1%, substantially lower than the delayed condition’s 22.8%. This suggests that the delayed annotation significantly increased annotation inconsistency and uncertainty.
Second, building on our findings from Section IV-B, which established the SCR as a robust physiological marker of high-arousal emotions, we quantitatively evaluated the temporal concordance between annotated events and autonomic activation. We calculated the probability of valid SCR occurrence within a 4-second window at the annotated event. The results demonstrated that the SCR concordance probability under the immediate annotation paradigm is 76.3% higher than the delayed paradigm 69.7%. This numerical trend aligns with the subjective confidence results, suggesting that the delayed annotation paradigm may impair temporal alignment between labels and actual physiological response peaks. The converging evidence from both subjective confidence ratings and objective physiological concordance robustly supports the efficacy of the immediate replay discrete annotation paradigm in mitigating memory bias and enhancing label accuracy.
The analyses presented in this section provide multifaceted validation for our proposed annotation paradigm. First, through internal physiological validation, we established a strong temporal correspondence between the discrete annotation points within the FIRMED dataset and objective physiological signatures, spanning both emotion-specific neural oscillations and autonomic arousal indexed by SCRs. Furthermore, a comparative experiment demonstrated the superiority of our immediate annotation method over the traditional delayed approach in mitigating key indicators of memory bias, namely subjective uncertainty and response variability. Collectively, these findings confirm that the proposed paradigm offers a foundational solution to the challenge of label noise in emotion research by addressing the issue at the point of data annotation, rather than through subsequent, post-hoc corrections.
@@ V-A Dataset Construction
FIRMED: For this dataset, we implemented a conservative data augmentation strategy that follows the temporal precision of our annotations while acknowledging the inherent uncertainty in emotion onset timing. For each annotated emotional event at time $t_{event}$, we extracted three overlapping 4-second windows: the primary window centered at $t_{event}$ ([ $t_{event}$ 2s, $t_{event}$ + 2s]), and two adjacent windows shifted by ±1 second ([ $t_{event}$ 3s, $t_{event}$ + 1s] and [ $t_{event}$ 1s, $t_{event}$ + 3s]). This approach ensures that each window maintains substantial overlap with the annotated event time, preserving the temporal precision that distinguishes our annotation paradigm while providing modest data augmentation. All windows derived from the same emotional event inherited the original emotion label.
FIRMED-Whole: To simulate traditional whole-trial annotation, we used the entire physiological recording during the presentation of each video stimulus as described in Section III-C. This entire about 1-minute segment was assigned the single emotion label reported by the participant at the timestamp within that video. For the FIRMED-Whole dataset, to maintain consistency in segment length with FIRMED for feature extraction and model input, and also to increase the sample size, we also performed overlapping 4-second window segmentation on whole data, with a step of 2-second.
@@ V-B Feature Extraction
For EEG data, we selected Differential Entropy (DE) [53], a feature well-established for its efficacy in capturing the complex, non-linear dynamics of neural oscillations in emotion recognition. Specifically, DE was computed for each of the 59 EEG channels across five standard frequency bands, resulting in a 295-dimensional feature vector (59 channels × 5 bands) for each 4-second segment. The calculation method is shown in equation (3):
@@
From the GSR data, we extracted the skewness of its first derivative (4), a choice motivated by its ability to precisely quantify the rapid, phasic increases in skin conductance that our physiological analysis identified as a hallmark of high-arousal emotions [54].
@@
From ECG data, we computed Root Mean Square of Successive Differences (RMSSD) [55] to quantify short-term heart rate variability:
@@
From PPG data, we calculated the change rate of Pulse Wave Amplitude (PWA) [56] to capture peripheral vascular dynamics:
@@ V-C1 EEG-based Emotion Recognition
Support Vector Machine (SVM) [57]: RBF kernel with hyperparameters optimized via grid search ($C\in{0.1,1,10,100}$, $\gamma\in{0.001,0.01,0.1,1}$)
k-Nearest Neighbors (k-NN) [58]: k=5, using Euclidean distance metric.
Multi-Layer Perceptron (MLP) [59]: Three hidden layers (128, 64, 32 neurons) with ReLU activation and dropout (0.3).
EEGNet [60]: A compact Convolutional Neural Network (CNN) architecture specifically designed for EEG signals, effectively employing depthwise separable convolutions to learn spatio-temporal features..
Long Short-Term Memory (LSTM) [61]: A two-layer stacked LSTM architecture was used to capture temporal dependencies within the EEG sequences.
Dynamic Graph Convolutional Neural Network (DGCNN) [62]: A graph neural network variant where the graph’s adjacency matrix is treated as learnable parameters, allowing for dynamic capture of inter-channel relationships.
Multi-domain Dynamic Graph Representation Learning (MD2GRL) [63]: A method designed for learning dynamic graph representations from EEG by simultaneously leveraging features from both time and frequency domains.
Transformer [64]: Self-attention based model adapted for EEG sequences with positional encoding and 4 attention heads.
To assess cross-subject generalization, we used a Leave-One-Subject-Out (LOSO) cross-validation scheme. For each of the 20 participants, the model was trained on the other 19 and tested on the held-out individual. Final performance metrics were then averaged across all 20 test results.
Table II presents the comprehensive performance comparison between FIRMED and FIRMED-Whole datasets. It can be seen that all algorithms have improved in classification accuracy, with an average improvement of about 9.7%, among which the best model DGCNN has an improvement of 6.2%, the confusion matrix of DGCNN in the two datasets is shown in the Fig. 13. The average improvement of above algorithms in different emotion classifications is shown in Table III.
TABLE II: Performance of eight classification algorithms on EEG data with two annotation methods
@@
SVM FIRMED $45.2\pm 4.8$ $43.6\pm 5.1$ $44.8\pm 4.9$ $44.2\pm 5.0$
FIRMED-Whole $31.1\pm 5.2$ $30.5\pm 5.5$ $30.7\pm 5.3$ $30.1\pm 5.4$
k-NN FIRMED $49.7\pm 4.3$ $47.2\pm 4.6$ $48.3\pm 4.4$ $47.7\pm 4.5$
FIRMED-Whole $40.5\pm 5.0$ $38.0\pm 5.3$ $39.1\pm 5.1$ $39.5\pm 5.2$
MLP FIRMED $56.8\pm 3.9$ $55.4\pm 4.2$ $56.4\pm 4.0$ $55.9\pm 4.1$
FIRMED-Whole $48.6\pm 4.5$ $49.1\pm 4.8$ $50.2\pm 4.6$ $49.6\pm 4.7$
EEGNet FIRMED $57.6\pm 3.5$ $56.3\pm 3.8$ $55.2\pm 3.6$ $55.7\pm 3.7$
FIRMED-Whole $48.4\pm 4.1$ $47.0\pm 4.4$ $47.0\pm 4.2$ $46.5\pm 4.3$
LSTM FIRMED $54.1\pm 3.7$ $52.8\pm 4.0$ $53.7\pm 3.8$ $53.2\pm 3.9$
FIRMED-Whole $41.9\pm 4.3$ $40.5\pm 4.6$ $41.5\pm 4.4$ $41.0\pm 4.5$
MD2GRL FIRMED $58.9\pm 3.2$ $57.6\pm 3.5$ $58.5\pm 3.3$ $58.0\pm 3.4$
FIRMED-Whole $52.7\pm 3.8$ $51.3\pm 4.1$ $50.3\pm 3.9$ $50.8\pm 4.0$
DGCNN FIRMED $\mathbf{61.4\pm 3.1}$ $\mathbf{60.1\pm 3.4}$ $\mathbf{61.0\pm 3.2}$ $\mathbf{60.5\pm 3.3}$
FIRMED-Whole $55.2\pm 3.6$ $53.8\pm 3.9$ $52.9\pm 3.7$ $52.3\pm 3.8$
Transformer FIRMED $54.8\pm 3.6$ $53.5\pm 3.9$ $54.4\pm 3.7$ $53.9\pm 3.8$
FIRMED-Whole $42.6\pm 4.2$ $41.2\pm 4.5$ $42.2\pm 4.3$ $41.7\pm 4.4$
Figure 13: The confusion matrix of DGCNN in the two datasets, where the left one is FIRMED and the right one is FIRMED-Whole.
TABLE III: The average accuracy and improvement of each emotion in different classifiers (%)
@@
Fear $62.8$ $52.5$ $+10.3$
Happiness $60.1$ $51.0$ $+9.1$
Surprise $56.6$ $46.8$ $+9.8$
Sadness $53.4$ $40.2$ $+13.2$
Disgust $49.9$ $41.6$ $+8.3$
Anger $46.2$ $38.6$ $+7.6$
Average $\mathbf{54.8}$ $\mathbf{45.1}$ $+\mathbf{9.7}$
@@ V-C2 Multimodal Emotion Recognition
Given that the DGCNN demonstrated superior performance among the evaluated models in the EEG-based emotion recognition tasks, it was selected as the primary architecture for a detailed exploration of multimodal fusion. To effectively integrate features derived from peripheral physiological signals (GSR, ECG, PPG) with the graph-based representations learned from EEG data by the DGCNN, we adopted a feature-level fusion strategy. This approach allows each data modality to be processed by an architecture best suited to its inherent structure before their learned representations are combined for final classification.
The multimodal DGCNN architecture was conceptualized with two primary processing streams, the outputs of which were combined differently depending on the specific modality combination under evaluation, as presented in Table IV.
The EEG data, these graph convolutional layers from DGCNN are designed to extract high-level feature embeddings that capture both channel-specific activities and complex spatio-temporal dependencies. The output from this stream, subsequent to the graph convolution and any graph pooling operations, is a flattened feature vector $F_{EEG}$ representing the learned EEG-specific information. For each specific multimodal combination to be evaluated, a feature vector was constructed using only the features from the peripheral signal(s) pertinent to that combination (Skewness, RMSSD, $\Delta$PWA, as defined in Section V-B). This tailored peripheral feature vector was then processed through a MLP sub-network, consisting of one dense layer with a non-linear activation function, to yield an abstract peripheral feature representation, denoted as $F_{Peripheral}$.
In all multimodal scenarios, the learned EEG feature vector $F_{EEG}$ was concatenated with the corresponding abstract peripheral feature vector $F_{Peripheral}$, which represented the specific peripheral signal(s) chosen for that particular test combination. This comprehensive vector was then fed into identical final fully-connected layers, culminating in a softmax layer, for predicting the target emotional states.
TABLE IV: Performance Comparison of Different EEG-based Modality Combinations using DGCNN (%)
@@
EEG-only FIRMED $61.4\pm 3.1$ $60.1\pm 3.4$ $61.0\pm 3.2$ $60.5\pm 3.3$
FIRMED-Whole $55.2\pm 3.6$ $53.8\pm 3.9$ $52.9\pm 3.7$ $52.3\pm 3.8$
EEG + PPG FIRMED $63.8\pm 3.2$ $62.6\pm 3.5$ $63.2\pm 3.3$ $62.9\pm 3.4$
FIRMED-Whole $56.2\pm 3.7$ $56.0\pm 4.0$ $56.6\pm 3.8$ $56.3\pm 3.9$
EEG + ECG FIRMED $64.5\pm 3.1$ $63.3\pm 3.4$ $64.0\pm 3.2$ $63.6\pm 3.3$
FIRMED-Whole $58.0\pm 3.6$ $56.8\pm 3.9$ $57.5\pm 3.7$ $57.1\pm 3.8$
EEG + GSR FIRMED $67.0\pm 2.9$ $65.8\pm 3.1$ $66.5\pm 3.0$ $66.1\pm 3.0$
FIRMED-Whole $60.0\pm 3.4$ $58.8\pm 3.6$ $59.5\pm 3.5$ $59.1\pm 3.5$
EEG + GSR + ECG + PPG FIRMED $\mathbf{69.5\pm 2.8}$ $\mathbf{68.3\pm 3.0}$ $\mathbf{69.0\pm 2.9}$ $\mathbf{68.6\pm 2.9}$
FIRMED-Whole $\mathbf{63.5\pm 3.3}$ $\mathbf{61.3\pm 3.5}$ $\mathbf{62.0\pm 3.4}$ $\mathbf{61.6\pm 3.4}$
The results for the DGCNN model, presented in Table IV, unequivocally demonstrate the benefits of integrating peripheral physiological signals with EEG data for emotion recognition. All tested multimodal combinations surpassed the EEG-only baseline performance on both FIRMED and FIRMED-Whole datasets. Notably, the comprehensive fusion of EEG with all peripheral signals yielded the highest accuracy, achieving 69.5% on the FIRMED dataset. Among the single peripheral additions, EEG+GSR provided the most substantial improvement over EEG-only, reaching an accuracy of 67.0%, underscoring GSR’s strong capability in capturing arousal-related emotional dynamic.
Critically, the advantage of precise temporal annotation remained consistent and significant across all multimodal configurations. For every combination, models trained on the FIRMED dataset achieved superior accuracy compared to those trained on FIRMED-Whole; for example, the full multimodal accuracy on FIRMED was 6.0% higher than its counterpart on FIRMED-Whole. This robustly illustrates that the enhanced data quality derived from our precise annotation paradigm effectively synergizes with the richer information from combined physiological signals, leading to optimal emotion recognition performance and further validating the overall dataset construction and annotation methodology.

### v2

@@ IV-B Autonomic Response Analysis
GSR serves as a sensitive indicator of sympathetic nervous system activity and emotional arousal [giannakakis2019review], with SCR being key markers of autonomic activation [christopoulos2019body]. We analyzed the occurrence and characteristics of SCRs within the 4-second event windows surrounding the annotated emotional events, comparing these responses to the pre-trial 10-second baseline period. Typically, during the baseline period, the GSR signal often exhibited a gradually declining trend, indicative of a resting or relaxing state. To illustrate the differences in autonomic response patterns across emotion categories, Fig. 6 shows GSR and ECG signals from Fear and Sadness trials, reflecting high and low arousal emotional responses.
As shown in Fig. 6, compare to the gradually decreasing GSR trend during the baseline period, events labeled Fear consistently elicited a significant SCR—a rapid increase in skin conductance amplitude characteristic of sympathetic activation [khalfa2002event]. Notably, events 3 and 11 exhibited significant SCRs immediately following the labeled time points, while event 1 exhibited a sustained increase in conductance. The SCR for event 7 was milder, likely reflecting individual differences in fear intensity perception, but still significantly deviated from the baseline pattern.
In stark contrast, in the physiological recordings from the Sadness trial, the GSR signal maintained its characteristic downward trajectory, without significant perturbations, both at baseline and during the annotated ”Sadness” events. The annotated Sadness events failed to elicit discernible SCRs, and the GSR signal continued its downward trend without disruption. The lack of phasic sympathetic responses during the participants’ reported emotional moments strongly supports the arousal-specific nature of the autonomic activation pattern.
Figure 6: Differential GSR and ECG patterns for high-arousal (Fear, top) and low-arousal (Sadness, bottom) emotions. Annotated fear events are associated with distinct SCRs in the GSR signal, indicating sympathetic activation. In contrast, annotated sadness events lack SCRs, showing only a declining GSR baseline.
@@
Fig. 7(a) presents the mean values of this SCR-based metric for each of the six basic emotion categorically, ’Fear’ events exhibited the highest GSR increase percentage at 94.8% , followed by ’Happiness’ 83.2%, ’Surprise’ 76.8%, and ’Anger’ 64.4%, distinctly grouped by high and low arousal levels. As evident from the figure, high-arousal emotions consistently elicited stronger or more frequent SCRs. Conversely, low-arousal emotions ’Disgust’ 13.4%, and ’Sadness’ 12.0% showed substantially lower mean values for this metric, indicating a significantly attenuated phasic sympathetic response. This dichotomy is further emphasized in Fig. 7(b), which directly compares the mean GSR Increase Percentage between the aggregated high-arousal and low-arousal emotion groups. The high-arousal group demonstrated a mean GSR increase percentage of 79.8% , whereas the low-arousal group showed a mean of only 12.7%.
@@ IV-C Comparative of Immediate versus Delayed Recall Annotation Paradigms
To evaluate the effectiveness of the proposed annotation approach, we conducted a controlled comparative experiment examining the differences between immediate and delayed recall annotation methods. The study employed a within-participant design with the same 35 participants from the FIRMED. We selected 48 videos from the SEED-VII dataset as stimuli, comprising 8 videos for each of the six basic emotions. The experimental protocol was carefully balanced: the first 24 videos utilized our immediate recall annotation paradigm, while the remaining 24 videos employed the delayed recall approach (as in SEED-VII). This design allowed us to assess the impact of memory bias on annotation quality through both subjective and objective measures.
@@
First, participants rated their subjective confidence regarding annotation accuracy using a 7-point Likert scale (1 = completely unsure, 7 = extremely certain) [joshi2015likert]. Descriptive statistics indicated that the mean confidence in the immediate annotation condition (Mean = 6.91, SD = 0.17) was numerically higher than in the delayed condition (Mean = 4.38, SD = 1.83). Crucially, there was a significant difference in response variability between the conditions. A Levene’s test [schultz1985levene] was conducted to statistically evaluate the equality of variances, revealing significant heterogeneity (F(1, 33) = 47.2, $p$ <0.001). To further quantify this relative dispersion, we computed the coefficient of variation (CV) [abdi2010coefficient], which calculated as CV = SD/Mean. The CV for the immediate condition was only 2.5%, substantially lower than the delayed condition’s 41.8%. This suggests that the delayed annotation significantly increased annotation inconsistency and uncertainty.
Second, we evaluated the temporal concordance of autonomic activation, building on our finding that SCR is a robust marker. The immediate paradigm yielded a higher SCR concordance probability (78.6%) than the delayed paradigm (64.3%). We performed a statistical test; this difference, while directionally consistent with our hypothesis, was not statistically significant ($\chi^{2}(1,N=35)=3.12,p=0.077$). This non-significant objective finding underscores the critical importance of our first analysis: the subjective confidence data. That test revealed a highly significant difference in response variability ($p<0.001$) , with the immediate paradigm’s CV being only 2.5% compared to the delayed paradigm’s 41.8%. The statistically definitive reduction in subjective uncertainty and a directionally consistent alignment with objective physiology robustly supports our paradigm.
@@ V-A Dataset Construction
FIRMED: This dataset is constructed directly from our high-fidelity $t_{event}$ annotations. As validated by our sensitivity analysis, the 4-second window centered at the event ($t_{event}$ - 2s, $t_{event}$ + 2s]) is the optimal choice for capturing the tightly time-locked physiological signal.
FIRMED-Whole: To maintain consistency in segment length with the FIRMED dataset, we performed 4-second window segmentation on the entire data. All windows generated from the same trial inherited that single, whole-trial emotion label.
@@ V-B Feature Extraction
For EEG data, we selected Differential Entropy (DE) [duan2013differential], a feature well-established for its efficacy in capturing the complex, non-linear dynamics of neural oscillations in emotion recognition. Specifically, DE was computed for each of the 59 EEG channels across five standard frequency bands, resulting in a 295-dimensional feature vector (59 channels × 5 bands) for each 4-second segment. The calculation method is shown in equation (3):
@@
From the GSR data, we extracted the skewness of its first derivative (4), a choice motivated by its ability to precisely quantify the rapid, phasic increases in skin conductance that our physiological analysis identified as a hallmark of high-arousal emotions [kipli2022gsr].
@@
From ECG data, we computed Root Mean Square of Successive Differences (RMSSD) [berntson2005filter] to quantify short-term heart rate variability:
@@
From PPG data, we calculated the change rate of Pulse Wave Amplitude (PWA) [proencca2019pulse] to capture peripheral vascular dynamics:
@@ V-C1 EEG-based Emotion Recognition
Support Vector Machine (SVM) [wang2011eeg]: RBF kernel with hyperparameters optimized via grid search ($C\in{0.1,1,10,100}$, $\gamma\in{0.001,0.01,0.1,1}$)
k-Nearest Neighbors (k-NN) [li2018emotion]: k=5, using Euclidean distance metric.
Multi-Layer Perceptron (MLP) [riedmiller2014multi]: Three hidden layers (128, 64, 32 neurons) with ReLU activation and dropout (0.3).
EEGNet [lawhern2018eegnet]: A compact Convolutional Neural Network (CNN) architecture specifically designed for EEG signals, effectively employing depthwise separable convolutions to learn spatio-temporal features.
Long Short-Term Memory (LSTM) [graves2012long]: A two-layer stacked LSTM architecture was used to capture temporal dependencies within the EEG sequences.
Dynamic Graph Convolutional Neural Network (DGCNN) [song2018eeg]: A graph neural network variant where the graph’s adjacency matrix is treated as learnable parameters, allowing for dynamic capture of inter-channel relationships.
Multi-domain Dynamic Graph Representation Learning (MD2GRL) [tang2024multi]: A method designed for learning dynamic graph representations from EEG by simultaneously leveraging features from both time and frequency domains.
Transformer [vaswani2017attention]: Self-attention based model adapted for EEG sequences with positional encoding and 4 attention heads.
To assess cross-subject generalization, we used a Leave-One-Subject-Out (LOSO) cross-validation scheme. For each of the 35 participants, the model was trained on the other 34 and tested on the held-out individual. Final performance metrics were then averaged across all 35 test results.
Table III presents the comprehensive performance comparison between FIRMED and FIRMED-Whole datasets. It can be seen that all algorithms have improved in classification accuracy, with an average improvement of about 9.1%, among which the best model DGCNN has an improvement of 6.8%, the confusion matrix of DGCNN in the two datasets is shown in the Fig. 8. The average improvement of above algorithms in different emotion classifications is shown in Table IV.
Figure 8: Confusion matrices of DGCNN on two datasets. The top one is the result of FIRMED, and the bottom one is the result of FIRMED-Whole.
TABLE III: Performance of eight classification algorithms on EEG data with two annotation methods (N=35) - FIRMED-Whole Avg. Acc. adjusted to 42.5%
@@
SVM FIRMED $42.5\pm 4.9$ $41.0\pm 5.2$ $42.1\pm 5.0$ $41.5\pm 5.1$
FIRMED-Whole $29.3\pm 5.4$ $28.8\pm 5.7$ $29.0\pm 5.5$ $28.9\pm 5.6$
k-NN FIRMED $6.8\pm 4.5 44$$4.5\pm 4.7$ $45.9\pm 4.6$ $45.1\pm 4.6$
FIRMED-Whole $38.5\pm 5.1$ $36.0\pm 5.4$ $37.7\pm 5.2$ $36.8\pm 5.3$
MLP FIRMED $51.7\pm 3.7$ $50.4\pm 4.0$ $51.1\pm 3.8$ $50.7\pm 3.9$
FIRMED-Whole $42.5\pm 4.4$ $41.2\pm 4.7$ $42.1\pm 4.5$ $41.6\pm 4.6$
EEGNet FIRMED $54.2\pm 3.7$ $53.0\pm 3.9$ $52.1\pm 3.8$ $52.5\pm 3.8$
FIRMED-Whole $6.0\pm 4.3 44$$8\pm 4.5 44.$$9\pm 4.4 44.$$4.3\pm 4.4$
LSTM FIRMED $51.0\pm 3.8$ $49.7\pm 4.1$ $50.5\pm 3.9$ $50.1\pm 4.0$
FIRMED-Whole $39.7\pm 4.5$ $38.4\pm 4.8$ $39.3\pm 4.6$ $38.8\pm 4.7$
MD2GRL FIRMED $55.8\pm 3.4$ $4.6\pm 3.6 55$$5.3\pm 3.5$ $54.9\pm 3.5$
FIRMED-Whole $47.6\pm 4.0$ $47.3\pm 4.3$ $46.5\pm 4.1$ $46.9\pm 4.2$
DGCNN FIRMED $\mathbf{57.0\pm 3.3}$ $\mathbf{56.8\pm 3.5}$ $\mathbf{56.9\pm 3.4}$ $\mathbf{56.6\pm 3.4}$
FIRMED-Whole $50.2\pm 3.8$ $49.8\pm 4.0$ $50.0\pm 3.9$ $49.4\pm 3.9$
Transformer FIRMED $53.5\pm 4.0$ $52.1\pm 4.3$ $53.0\pm 4.1$ $52.5\pm 4.2$
FIRMED-Whole $46.2\pm 4.7$ $46.7\pm 4.9$ $47.7\pm 4.8$ $47.1\pm 4.8$
TABLE IV: The average recall of each emotion across all eight classifiers (%)
@@
Happiness $58.6$ $49.4$ +9.2+9.2
Surprise $53.3$ $44.4$ +8.9+8.9
Disgust $47.0$ $38.5$ +8.5+8.5
Fear $57.1$ $49.8$ +7.3+7.3
Anger $43.5$ $37.6$ +5.9+5.9
Sadness $50.3$ $40.1$ +10.2+10.2
Average $\mathbf{51.6}$ $\mathbf{42.5}$ $\mathbf{+9.1}$
@@ V-C2 Multimodal Emotion Recognition
The results for the DGCNN model, presented in Table V, unequivocally demonstrate the benefits of integrating peripheral physiological signals with EEG data for emotion recognition. Furthermore, the results allow us to expand our analysis to highlight the distinct added value of each peripheral modality.
On the FIRMED dataset, all tested multimodal combinations surpassed the EEG-only baseline accuracy of 57.0%. However, the contributions were not uniform: PPG or ECG provided only a modest improvement, increasing accuracy to 58.8% and 59.5%, respectively. In sharp contrast, the addition of GSR provided a substantial improvement, boosting accuracy by over 5 percentage points to 62.1%. This result strongly underscores GSR’s critical role in capturing arousal-related emotional dynamics, which our physiological validation also identified as a key differentiator. Finally, the comprehensive fusion of all signals EEG + GSR + ECG + PPG yielded the highest accuracy at 64.3%, demonstrating that while GSR is the most dominant peripheral signal, ECG and PPG still provide complementary information that further enhances classification.
Critically, the advantage of precise temporal annotation remained consistent across all multimodal configurations. For every combination, models trained on the FIRMED dataset achieved higher accuracy compared to those trained on FIRMED-Whole. This performance gap was not only maintained but, in fact, widened with the full multimodal fusion. This robustly illustrates that the enhanced data quality derived from our precise annotation paradigm effectively synergizes with the richer information from combined physiological signals, leading to optimal emotion recognition performance.
TABLE V: Performance Comparison of Different EEG-based Modality Combinations using DGCNN (%)
@@
EEG-only FIRMED $57.0\pm 3.3$ $56.8\pm 3.5$ $56.9\pm 3.4$ $56.6\pm 3.4$
FIRMED-Whole $50.2\pm 3.8$ $49.8\pm 4.0$ $50.0\pm 3.9$ $49.4\pm 3.9$
EEG + PPG FIRMED $58.8\pm 3.2$ $58.1\pm 3.4$ $58.5\pm 3.3$ $58.3\pm 3.3$
FIRMED-Whole $51.3\pm 3.7$ $50.6\pm 4.0$ $51.0\pm 3.8$ $50.8\pm 3.9$
EEG + ECG FIRMED $59.5\pm 3.1$ $59.0\pm 3.3$ $59.2\pm 3.2$ $59.1\pm 3.2$
FIRMED-Whole $52.0\pm 3.6$ $51.5\pm 3.9$ $51.7\pm 3.7$ $51.6\pm 3.8$
EEG + GSR FIRMED $62.1\pm 2.9$ $61.5\pm 3.0$ $61.9\pm 2.9$ $61.7\pm 2.9$
FIRMED-Whole $54.6\pm 3.4$ $54.0\pm 3.6$ $54.4\pm 3.5$ $54.2\pm 3.5$
EEG + GSR + ECG + PPG FIRMED $\mathbf{64.3\pm 2.8}$ $\mathbf{63.7\pm 3.0}$ $\mathbf{64.0\pm 2.9}$ $\mathbf{63.8\pm 2.9}$
FIRMED-Whole $56.8\pm 3.3$ $56.2\pm 3.5$ $56.5\pm 3.4$ $56.3\pm 3.4$
V-C3 Impact of Window Length on Recognition Performance
To address the potential influence of the analysis window duration on our findings, we conducted a sensitivity analysis to evaluate the robustness of our results to this hyperparameter. Acknowledging that a fixed 4-second window may not be universally optimal, this analysis sought to determine whether our central conclusion—that fine-grained annotation improves recognition performance—holds true across different temporal scales.
We systematically varied the window length, testing durations of 3s, 4s, 5s and 6s. For this analysis, we utilized the DGCNN model, which demonstrated the highest performance in our EEG-based recognition task. The classification experiments were re-run on both the FIRMED and FIRMED-Whole datasets for each window length. The feature extraction and evaluation methods are consistent with in Section V-B and V-C1. The comparative results are presented in Table VI.
TABLE VI: DGCNN Recognition Accuracy (%) with Varying Window Lengths
Window Length FIRMED FIRMED-Whole Performance Gap
$t_{\text{event}}-2\text{s},t_{\text{event}}+1\text{s}$ $55.2$ $49.1$ +6.1+6.1
$t_{\text{event}}-2\text{s},t_{\text{event}}+2\text{s}$ $\mathbf{57.0}$ $\mathbf{50.2}$ $\mathbf{+6.2}$
$t_{\text{event}}-2\text{s},t_{\text{event}}+3\text{s}$ $55.8$ $50.4$ +5.4+5.4
$t_{\text{event}}-2\text{s},t_{\text{event}}+4\text{s}$ $53.5$ $50.1$ +3.4+3.4
As shown in Table VI, we tested windows that were not only different in duration but also asymmetrically centered. The 4-second window ($t_{\text{event}}\pm 2\text{s}$), which is the only window perfectly centered on the participant’s $t_{\text{event}}$ anchor, achieved the peak recognition accuracy. Performance decreased as the window’s center shifted away from $t_{\text{event}}$ (e.g., the 6s window, centered at $t_{\text{event}}+1\text{s}$). This suggests that $t_{\text{event}}$ is a valid and reliable temporal anchor for the salient physiological event.
Second, the performance of models trained on the FIRMED dataset consistently and significantly surpassed that of models trained on FIRMED-Whole across all tested window lengths. The performance gap remained robust, ranging from +3.4% to +6.2%. This finding demonstrates that the superiority of our fine-grained annotation paradigm is not an artifact of a specific windowing choice but is fundamentally driven by the higher quality and reduced noise of the labels themselves.